The prescribed dosage of a medicine varies directly with weight. If the dosage for a 120 pound person is 8 ml, what should the dosage be for a person weighing 210 pounds?
Answer by richwmiller(17219)

y=kx
8=120k
8/120=k
1/15=k
y=1/15*x
y=1/15*210
y=(1/15)*3*5*7*2
y=7*2=14 mL